OpenStudy (anonymous):

Updated Batting Averages are computed using the formula: A= (B+K) over G •A = batting average •B = times at bat multiplied by player’s seasonal average •K = hits in the new game •G = times at bat plus times at bat in a new game solve A= (B + K) over G Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Round your answer to the nearest whole number.

OpenStudy (anonymous):

a = (b + k) / g Multiply both sides by g: a*g = b + k Subtract b from both sides (a*g) - b = k For George, k = (0.319)(46) - 10.7 = 3.974 = 4 to the nearest whole number. This is just arithmetic now. You can do it as well as I.

OpenStudy (anonymous):